OpenStudy (anonymous):

23=0.5xb This equation could be used to answer which question?

OpenStudy (apoorvk):

This question!

OpenStudy (anonymous):

a) 23 is 5% of what number

OpenStudy (anonymous):

xb = 23*2 = 46 It might be used for finding the area of rectangle...

OpenStudy (anonymous):

b) what is 23% of 0.5

OpenStudy (apoorvk):

ohkay, nope. 0.5% = 0.005 actually. so check it out!

OpenStudy (anonymous):

c)50% of what number is 23

OpenStudy (anonymous):

5x/100 = 23 x = (23*100)/5 = 460

OpenStudy (anonymous):

what is 50% of 23

OpenStudy (apoorvk):

@waterineyes I think that 'x' represents a "multiplication' sign here, not 'eks'. Comments? -.-

OpenStudy (anonymous):

yes x does

OpenStudy (apoorvk):

50% = 0.5 any idea now?

OpenStudy (anonymous):

it's b ??

OpenStudy (anonymous):

b) (23*0.5)/100 = 23/200 = 11.5/100 = 0.115

OpenStudy (anonymous):

so it is b

OpenStudy (anonymous):

@apoorvk , I am writing x there not multiplication and that is the answer of a). I always sign multiply as *..

OpenStudy (anonymous):

so it's a

OpenStudy (apoorvk):

Umm, not really. *Please try again* PS - *Think and try again*

OpenStudy (anonymous):

C) 50x/100 = 23 x = 46 Therefore I think C is the answer..

OpenStudy (apoorvk):

Oui^ - should have let her race her brain first though -_-

OpenStudy (anonymous):

Sorry apoorvk..

OpenStudy (apoorvk):

Lol nothing to be sorry about aye. Good job aye aye! xD